OpenStudy (anonymous):

A $200.00 sweater has been marked down to $120.00. What percentage discount from the original price does the markdown represent? 50% off 30% off 20% off 60% off 40% off

OpenStudy (anonymous):

(200-120)/200 = 8/20 = 2/5 = 40%

OpenStudy (anonymous):

No matter what way i look at math problems i just dont get them